BAK Corp. is considering purchasing one of two new diagnostic machines. Either machine would make it possible for the company to bid on jobs that it currently isn’t equipped to do. Estimates regarding each machine are provided below.

Machine A Machine B
Original cost $74,000 $179,000
Estimated life 8 years 8 years
Salvage value 0 0
Estimated annual cash inflows $19,500 $39,500
Estimated annual cash outflows $4,800 $9,800



Click here to view PV table.

Calculate the net present value and profitability index of each machine. Assume a 9% discount rate. (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round answer for present value to 0 decimal places, e.g. 125 and profitability index to 2 decimal places, e.g. 10.50. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Machine A Machine B
Net present value enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Profitability index enter the profitability index rounded to 2 decimal places enter the profitability index rounded to 2 decimal places


Which machine should be purchased?

select a machine                                                          Machine AMachine B should be purchased.
Answer #1

Calculate following

Machine A Machine B
Net present value (14700*5.53482)-74000 = 7362 (29700*5.53482)-179000 = -14616
Profitability index 81362/74000 = 1.10 164384/179000 = 0.92

Machine A should be purchased